# Unveiling the Mathematical Reasoning in DeepSeek Models: A Comparative Study of Large Language Models

### v1

@@ Abstract
With the rapid evolution of Artificial Intelligence (AI), Large Language Models (LLMs) have reshaped the frontiers of various fields, spanning healthcare, public health, engineering, science, agriculture, education, arts, humanities, and mathematical reasoning. Among these advancements, DeepSeek models have emerged as noteworthy contenders, demonstrating promising capabilities that set them apart from their peers. While previous studies have conducted comparative analyses of LLMs, few have delivered a comprehensive evaluation of mathematical reasoning across a broad spectrum of LLMs. In this work, we aim to bridge this gap by conducting an in-depth comparative study, focusing on the strengths and limitations of DeepSeek models in relation to their leading counterparts. In particular, our study systematically evaluates the mathematical reasoning performance of two DeepSeek models alongside five prominent LLMs across three independent benchmark datasets. The findings reveal several key insights: 1). DeepSeek-R1 consistently achieved the highest accuracy on two of the three datasets, demonstrating strong mathematical reasoning capabilities. 2). The distilled variant of LLMs significantly underperformed compared to its peers, highlighting potential drawbacks in using distillation techniques. 3). In terms of response time, Gemini 2.0 Flash demonstrated the fastest processing speed, outperforming other models in efficiency, which is a crucial factor for real-time applications. Beyond these quantitative assessments, we delve into how architecture, training, and optimization impact LLMs’ mathematical reasoning. Moreover, our study goes beyond mere performance comparison by identifying key areas for future advancements in LLM-driven mathematical reasoning. This research enhances our understanding of LLMs’ mathematical reasoning and lays the groundwork for future advancements.
@@ 1 Introduction
As Artificial Intelligence (AI) continues to advance at an unprecedented pace, an array of powerful Large Language Models (LLMs) has emerged, including OpenAI’s GPT-4o, o1, o3, Claude 3.5, Llama 3.2, Qwen 2.5, and Gemini 2.0 [1, 12, 28, 30, 48, 57, 69, 82, 102]. Driven by cutting-edge advancements in Deep Neural Networks (DNNs), these models integrate human cognition elements to enhance problem-solving and decision-making [49]. Serving as a guiding light in natural language processing (NLP) [65], healthcare [21, 60, 61], clinical textual processing [22, 59, 104], biomedical image analyses [60], code generation [58], decision support [51], multimodal data analytics [83, 90], and mathematical reasoning [2, 52, 72, 76, 91, 103, 107], LLMs push the boundaries of AI capabilities, approximating human-like reasoning through sophisticated statistical inference [1, 48, 69]. However, despite their transformative potential, these models face notable limitations. Their high computational demands pose significant barriers to broader accessibility, making large-scale implementation costly [69]. Moreover, while LLMs perform well in general contexts, they often struggle with specialized tasks, exhibiting inconsistencies in performance. Multimodal models, for example, continue to face challenges in spatial reasoning and real-world physics, while AI-assisted code generation frequently produces syntactically correct yet functionally flawed outputs, requiring human oversight [58, 103]. These constraints underscore the ongoing need for refinement and innovation in AI research to bridge the gap between artificial and human intelligence.
In particular, GPT-4o, released by OpenAI in May 2024, is a multimodal model capable of processing text, images, and voice with remarkable efficiency. Leveraging an advanced transformer architecture, it surpasses GPT-3 in critical areas such as mathematical reasoning and language comprehension [1, 23, 44, 75]. With an estimated 2 trillion parameters, GPT-4o is significantly larger than its predecessors, enabling substantial improvements in performance and adaptability. Meanwhile, other cutting-edge models, such as o1 and o3, have been introduced in 2024 and 2025, respectively. In particular, o1 enhances reasoning capabilities by incorporating a Chain-of-Thought (CoT) approach [80], where it renders intermediate reasoning steps before reaching a final answer closely mirroring the way humans process complex problems [103]. More significantly, o3 advances reasoning performance through a "simulated reasoning" process, in which the model actively generates and evaluates multiple solution paths [6]. By pausing, reflecting, and adjusting its approach before delivering a final answer, o3 exhibits human-like reasoning, allowing for more nuanced and adaptive problem-solving, particularly in highly complex scenarios [67]. However, none of the GPT models, including GPT-4o, o1, and o3, are open-source, as they remain proprietary.
@@
In addition, Claude 3.5, released in 2024, is built on previous versions. It emphasizes safety, alignment, and performance, with improvements in reasoning, language understanding, and handling complex tasks like text and code generation [30]. With 250B parameters, it surpasses earlier models in accuracy and ethical alignment. It supports up to 200K tokens for extended context, enabling better processing of larger inputs. Notably, enhanced by reinforcement learning from human feedback (RLHF) [41, 85] and Constitutional AI, it reduces undesirable responses, biases, and better aligns with human intent. Claude 3.5 excels in specialized areas like coding and scientific reasoning, with improved transparency and ethical safeguards [30, 69].
@@
Furthermore, Llama-3.3, presented in 2024, is the latest version of LLM in Meta AI family, following Llama-1 and Llama-2. Llama-3.3 advances further with 70B parameters and 128K token context window, improved by grouped-query attention for better efficiency [28, 63]. Llama 3.1 excels in coding, logical problem solving, and low-resource language tasks. Unlike closed models such as GPT series, it remains open-weight and freely accessible for research and commercial use, but is restricted to text-only input [26, 63, 69]. Safety measures, such as automated red-teaming and filtered training data, help minimize undesirable outputs. Moreover, Gemini 2.0 Flash is the latest multimodal LLM of Google, building on versions 1.0 and 1.5 to offer more robust generative AI capabilities across text, images, audio, and video.
@@
Moreover, Gemini 2.0 Flash, initially introduced as an experimental variant, provides significant speed and efficiency gains over its predecessor, Gemini 1.5 Flash, without sacrificing the efficiency [43, 45, 82]. Notably, it outperforms Gemini 1.5 Pro on key benchmarks while operating at twice the speed. It enables the incorporation of agentic AI and native use, allowing the model to call external functions (Google Search and Maps) and integrate streaming data for expanded real-time applications. By combining better performance in tasks such as math, code generation, and multilingual audio output with enhanced efficiency. Gemini 2.0 aims to deliver comprehensive, cost-effective AI solutions for both developers and end users [43, 82]. Lastly, Qwen2.5, released in September 2024, is the latest iteration in the Qwen series, following Qwen2 in June 2024 and the original Qwen in August 2023. Qwen1.5 featured models up to 72B parameters, emphasizing efficiency and open-source accessibility. Qwen2 introduced improved reasoning, multilingual support, and coding capabilities, with models scaling up to 72.71B parameters. [94, 106]
@@
Besides, established in 2023 as a research initiative to push the boundaries of artificial general intelligence (AGI), DeepSeek models set out to overcome existing limitations by developing specialized models focused on efficiency, adaptability, and domain expertise [8, 20, 38, 37, 57, 48, 56, 76]. In 2024, the Mixture-of-Experts (MoE), an efficiency-driven architecture that leverages sparse activation to reduce computational overhead [38, 56, 57], for DeepSeek was introduced. This was followed by the launch of DeepSeek Coder, a suite of code-focused models ranging from 1B to 33B parameters, designed to streamline software development workflows. Meanwhile, DeepSeek Math, trained on 120B math-related tokens, was developed to handle advanced mathematical and symbolic reasoning tasks [76]. Expanding its model portfolio, DeepSeek introduced the V2 and V3 series [56, 57]. V2 implemented multi-head latent attention (MLA) alongside a MoE system with 236B parameters in total, of which only 21B were active per query, optimizing computational efficiency [56]. V3, an open-source model, further enhanced efficiency with 671B total parameters, activating only 37B per query, excelling in complex reasoning tasks while minimizing resource demands and reliance on supervised data [57]. In 2025, DeepSeek made a significant breakthrough with R1 Zero, incorporating self-verification, reflection, and extended CoTs. Notably, DeepSeek-R1 was recently presented, specifically designed for mathematical, coding, and logical problem-solving, enhancing autonomous decision-making and precision in both research and enterprise applications [37, 38, 76]. To extend accessibility, an open-sourced suite of distilled models for DeepSeek, optimized for deployment in resource-constrained environments such as edge computing platforms and low-memory systems. These models preserve scalability and cost-effectiveness, making cutting-edge AI more accessible across diverse applications.
@@
Notably, mathematical reasoning usually poses intricate challenges without straightforward solutions [3, 55]. Unlike routine tasks that follow established frameworks, mathematical reasoning demands creativity, abstract thinking, and advanced cognitive skills. In this study, we systematically evaluate the mathematical reasoning performance of a wide array of LLMs, with a particular emphasis on DeepSeek models, to further unveil machine creativity in mathematical reasoning. Leveraging multiple independent datasets and quantitative metrics, we conduct a comprehensive empirical analysis to assess and compare their capabilities. Our findings provide a detailed evaluation of prominent LLMs in mathematical reasoning while highlighting the advancements and unique strengths of DeepSeek models.
2 Related Works
Recent research has increasingly focused on summarizing and evaluating the performance of LLMs in problem-solving and code generation, areas where general-purpose LLMs excel in text-based tasks but often struggle with mathematical precision and structured reasoning [24, 36, 48, 77, 86, 97]. To address these limitations, LLMs have prioritized enhancing reasoning capabilities and improving computational efficiency in next-generation models, aiming to bridge the gap between linguistic fluency and robust problem-solving skills.
Specifically, in 2024, Wang et al., summarized and detailed the specifications of Multimodal Large Language Models (MLLMs), standing out at the forefront of AI [83]. In brief, Wang et al., overviewed a wide array of data modalities, such as text, images, audio, and sequential data [83]. In fact, MLLMs play a crucial role in multimodal understanding tasks by integrating text and image information to achieve more intelligent and comprehensive understanding and reasoning. In this field, Wang et al., outlined the developments, advancements, and utilization of multiple MLLMs, including MiniGPT-4, InstructBLIP, and Wiki-LLaVA are three highly regarded models, along with other related MLLMs such as 3DMIT, GroundingGPT, ModaVerse, Vary-toy, LLaVAMOLE, and CogCom [83].
In addition, at the end of 2024, Zhong et al. established a comprehensive and extensive evaluation of o1-preview (the early version of o1) to showcase the advancement of OpenAI’s o1-preview across a diverse array of complex reasoning tasks, spanning multiple domains, including computer science, mathematics, natural sciences, medicine, linguistics, and social sciences [103]. Through extensive testing, o1-preview highlighted remarkable capabilities, often achieving human-level or superior performance in areas ranging from coding challenges to scientific reasoning and from language processing to creative problem-solving [103]. Key findings include: 1). 83.3% success rate in solving complex competitive programming problems, surpassing human experts. 2). Superior ability in generating coherent and accurate radiology reports, outperforming other evaluated models. 3). 100% accuracy in high school-level mathematical reasoning tasks, providing detailed step-by-step solutions. 4). Advanced natural language inference capabilities across general and specialized domains like medicine. 5). Impressive performance in chip design tasks, outperforming specialized models in areas such as script generation and bug analysis. 6). Remarkable proficiency in anthropology and geology, demonstrating deep understanding and reasoning in these specialized fields. 7). Comprehensive financial knowledge and strong capabilities in quantitative investing. 8). Effective performance in social media analysis, including sentiment analysis and emotion recognition. 9). Excellent performance in educational measurement and psychometrics, demonstrating a solid grasp of standard psychometric concepts that are equivalent to or beyond a first-year master’s or doctoral student’s level.
Furthermore, Yang et al. investigated the capabilities of advanced LLMs, particularly the o1 model, in literary analysis [96]. Given the prestige of the Nobel Prize and its emphasis on cultural, historical, and linguistic depth, applying LLMs to these works offers valuable insights into both human and AI approaches to literary interpretation [96]. The study employed qualitative and quantitative evaluations to assess coherence, creativity, and fidelity to the text, shedding light on the strengths and limitations of AI in domains traditionally dominated by human expertise. While LLMs demonstrated remarkable analytical capabilities, particularly in structured tasks, they struggled with emotional nuance and coherence—areas where human interpretation remains unparalleled. This research underscores the transformative potential of human-AI collaboration in the humanities, paving the way for new opportunities in literary studies, textual analysis, and interdisciplinary research. By leveraging AI’s analytical power alongside human intuition, this study highlights promising directions for AI-assisted literary interpretation and beyond.
Moreover, Xu et al. demonstrated that LLMs have made significant advancements in clinical decision-making, particularly those leveraging in-context demonstrations and specialized medical fine-tuning [92]. These models exhibit strong performance in medical language processing but continue to face challenges in real-time adaptability, multi-step reasoning, and handling complex medical tasks. Agent-based AI systems aim to overcome these limitations by incorporating reasoning traces, contextual tool selection, knowledge retrieval, and both short- and long-term memory. These features enable medical AI agents to manage complex clinical scenarios where decision-making requires real-time interaction with the environment [92]. Unlike conventional model-based approaches that treat medical queries as isolated questions, medical AI agents approach them as dynamic, multi-faceted tasks, allowing them to function more like human doctors. Xu also investigated the selection of the backbone LLM for medical AI agents, which serves as the foundation for their reasoning and action generation. Specifically, Xu et al. [92] examined the capabilities of the emerging o1 model and its impact on agents’ reasoning, tool-use adaptability, and real-time information retrieval across diverse clinical scenarios, including high-stakes environments such as intensive care units (ICUs). The findings highlighted o1’s potential to enhance diagnostic accuracy and consistency, paving the way for more intelligent, responsive AI tools that support improved patient outcomes and more effective decision-making in clinical practice [92].
Notably, Ahn et al. [3] addressed four critical aspects of LLMs in mathematical reasoning: 1). mathematical problem types and datasets, 2). techniques for enhancing LLM performance (such as prompt engineering and fine-tuning), 3). factors affecting model effectiveness (including scale and pre-training data), and 4). challenges such as brittleness and inconsistency, which often cause models to generate different answers for similar problems.
Besides, Liang et al. presented a comprehensive overview on the applications of AI in mathematical research, highlighting the transformative role AI has begun to play in this domain [55]. Traditionally, AI advancements have heavily relied on theoretical foundations provided by mathematics and statistics. However, recent developments in AI, particularly in reinforcement learning (RL) and LLMs, have demonstrated the potential for AI to contribute back to mathematics by offering flexible algorithmic frameworks and powerful inductive reasoning capabilities that support various aspects of mathematical research. This survey aimed to establish a bridge between AI and mathematics, providing insights into the mutual benefits and fostering deeper interdisciplinary understanding. In particular, Liang et al. discussed that while current AI and LLMs may struggle with complex deductive reasoning, their inherent creativity, the capability to generate outputs at high throughput based on recognition of shallow patterns, holds significant potential to support and inspire mathematical research [55]. Furthermore, Liang et al. [55] addressed the lack of cross-disciplinary communication: mathematicians may not fully comprehend the latest advances in AI, while AI researchers frequently prioritize benchmark performance over real-world applications in frontier mathematical research.
Lastly, this work seeks to close that gap, offering a detailed exploration of AI fundamentals, its strengths, and its emerging applications in the mathematical sciences. While various studies have begun exploring in-depth comparisons between DeepSeek-R1 and other leading models [37, 38, 76], few provide a comprehensive analysis of their mathematical reasoning capabilities. Thus, in this work, we introduce a holistic empirical study to systematically compare the mathematical reasoning performance of DeepSeek-R1 with its peer models. Our goal is to contribute to a deeper understanding of DeepSeek-R1’s strengths and limitations while inspiring further advancements in the future works, using a wide array of peer models, datasets, and quantitative metrics.
3 Methodology
@@
In this section, we provide a comprehensive overview of our experimental framework, detailing the datasets utilized, the quantitative evaluation metrics, as well as the overall methodology. We also present a justification for selecting these datasets as benchmarks, emphasizing their relevance, diversity, and suitability in effectively assessing LLM performance in mathematical reasoning. By establishing a consistent evaluation framework, we aim to ensure a thorough and insightful comparison of model capabilities across various reasoning tasks.
Figure 1: This figure illustrates the experimental framework of this work, including five vital components.
3.1 Experimental Frameworks
The framework of this empirical study is illustrated in Figure 1 and consists of five key components: 1). Benchmark Data: We utilized three representative benchmark datasets [16, 39, 62] to evaluate the mathematical reasoning capabilities of multiple LLMs. 2). Zero-Shot Prompting: Zero-shot prompting was employed to design all prompts [42], ensuring a standardized evaluation approach. 3). Language Model Selection: A diverse set of representative LLMs was selected, including Google’s Gemini 2.0 Flash [45], OpenAI’s GPT-4o [1], o1-mini [46], o1 [103], o3-mini [6], DeepSeek’s DeepSeek-R1-Distill-Qwen-1.5B [37], hereafter referred to as DeepSeek-1.5B, and DeepSeek-R1 [56]. 4). Evaluation: Accuracy was used as a critical quantitative metric to assess and compare the mathematical reasoning performance of these models [17, 18, 19]. 5). Comparison Between Models: A comprehensive statistical analysis was conducted to evaluate and contrast the performance of each LLM, offering deeper insights into their mathematical reasoning capabilities. Overall, this structured framework enabled a holistic and systematic comparison of LLMs in mathematical reasoning.

### v2

# Evaluating Mathematical Reasoning Across Large Language Models: A Fine-Grained Approach

@@ Abstract
With the rapid advancement of Artificial Intelligence (AI), Large Language Models (LLMs) have significantly impacted a wide array of domains, including healthcare, engineering, science, education, and mathematical reasoning. Among these, mathematical reasoning remains a particularly challenging capability, often requiring multi-step logic and abstract generalization. While prior work has explored LLM performance on reasoning tasks, comprehensive evaluations that span both depth and breadth across model families remain limited. In this study, we present a systematic evaluation of mathematical reasoning abilities across eight leading LLMs, including two recent DeepSeek models, using three independent benchmark datasets. Our analyses reveal several key findings: (1) DeepSeek-R1 performs competitively with o1 across most domains and achieves the highest accuracy on the MMLU Formal Logic benchmark; (2) distilled variants, such as DeepSeek-1.5B, exhibit substantial performance degradation; and (3) Gemini 2.0 Flash achieves the lowest response latency. Beyond quantitative metrics, we explore how architectural choices, training paradigms, and optimization strategies contribute to variation in reasoning performance. These findings provide new insights into the capabilities and limitations of current LLMs in mathematical domains, and offer guidance for the development of future models better aligned with rigorous reasoning demands.
@@ 1 Introduction
As Artificial Intelligence (AI) continues to advance at an unprecedented pace, an array of powerful Large Language Models (LLMs) has emerged, including OpenAI’s GPT-4o, o1, o3, Claude 3.5, Llama 4, Qwen 3, Gemini 2.0 and so on [Achiam et al., 2023, Chang et al., 2024, Dubey et al., 2024, Enis and Hopkins, 2024, Kasneci et al., 2023, Liu et al., 2024b, Team et al., 2024, Zhao et al., 2023]. Driven by cutting-edge advancements in Deep Neural Networks (DNNs), these models integrate human cognition elements to enhance problem-solving and decision-making [Lappin, 2024]. Serving as a guiding light in natural language processing (NLP) [Min et al., 2023], healthcare [Dai et al., 2023, Liu et al., 2023c, d], clinical textual processing [Dai et al., 2025, Liu et al., 2023b, Zhong et al., 2023], biomedical image analyses [Liu et al., 2023c], code generation [Liu et al., 2023a], decision support [Li et al., 2022], multimodal data analytics [Wang et al., 2024, Xiao et al., 2024], and mathematical reasoning [Ahn et al., 2024a, Li et al., 2024b, Peng et al., 2024, Shao et al., 2024, Xin et al., 2024, Zhong et al., 2024, Zhuang et al., 2024], LLMs push the boundaries of AI capabilities, approximating human-like reasoning through sophisticated statistical inference [Achiam et al., 2023, Kasneci et al., 2023, Neha and Bhati, 2025]. However, despite their transformative potential, these models face notable limitations. Their high computational demands pose significant barriers to broader accessibility, making large-scale implementation costly [Neha and Bhati, 2025]. Moreover, while LLMs perform well in general contexts, they often struggle with specialized tasks, exhibiting inconsistencies in performance. Multimodal models, for example, continue to face challenges in spatial reasoning and real-world physics, while AI-assisted code generation frequently produces syntactically correct yet functionally flawed outputs, requiring human oversight [Liu et al., 2023a, Zhong et al., 2024]. These constraints underscore the ongoing need for refinement and innovation in AI research to bridge the gap between artificial and human intelligence.
In 2024, OpenAI released GPT-4o, a multimodal model capable of processing text, images, and audio. With estimated 2 trillion parameters, it significantly surpasses GPT-3 in tasks such as reasoning, coding, and clinical diagnostics [Achiam et al., 2023, Dale, 2021, Hurst et al., 2024, Shahriar et al., 2024]. Meanwhile, other cutting-edge models, such as o1, o3 and recently o4-mini, have been introduced between 2024 and 2025. In particular, o1 enhances reasoning capabilities by incorporating a Chain-of-Thought (CoT) approach [Stechly et al., 2025, Zhong et al., 2024], where it renders intermediate reasoning steps before reaching a final answer closely mirroring the way humans process complex problems [Arrieta et al., 2025].
@@
In addition, Claude 3.5, released in 2024, emphasizes safety, alignment, and performance, with improvements in reasoning, language understanding, and handling complex tasks like text and code generation [Enis and Hopkins, 2024]. With 250B parameters, it surpasses earlier models in accuracy and ethical alignment. It supports up to 200K tokens for extended context. Notably, enhanced by reinforcement learning from human feedback (RLHF) [Hu et al., 2024, Wen et al., 2024] and Constitutional AI, it reduces undesirable responses, biases, and better aligns with human intent. Claude 3.5 excels in specialized areas like coding and scientific reasoning, with improved transparency and ethical safeguards [Enis and Hopkins, 2024, Neha and Bhati, 2025].
@@
Furthermore, Llama-3.3, presented in 2024, is the latest version of LLM in Meta AI family, following Llama-1 and Llama-2. Llama-3.3 advances further with 70B parameters and 128K token context window, improved by grouped-query attention for better efficiency [Dubey et al., 2024, Lu et al., 2024]. Llama 3.1 excels in coding, logical problem solving, and low-resource language tasks. Unlike closed models such as GPT series, it remains open-weight and freely accessible for research and commercial use, but is restricted to text-only input [Dominguez-Olmedo et al., 2025, Lu et al., 2024, Neha and Bhati, 2025].
@@
Moreover, Gemini 2.0 Flash, initially introduced as an experimental variant, provides significant speed and efficiency gains over its predecessor, Gemini 1.5 Flash, without sacrificing the efficiency [Huang and Zhang, 2024, Imran and Almusharraf, 2024, Team et al., 2024]. Notably, it outperforms Gemini 1.5 Pro on key benchmarks while operating at twice the speed. It enables the incorporation of agentic AI and native use, allowing the model to call external functions (Google Search and Maps) and integrate streaming data for expanded real-time applications. By combining better performance in tasks such as math, code generation, and multilingual audio output with enhanced efficiency. Gemini 2.0 aims to deliver comprehensive, cost-effective AI solutions for both developers and end users [Huang and Zhang, 2024, Team et al., 2024]. Lastly, Qwen2.5, released in September 2024, is the latest iteration in the Qwen series, following Qwen2 in June 2024 and the original Qwen in August 2023. Qwen1.5 featured models up to 72B parameters, emphasizing efficiency and open-source accessibility. Qwen2 introduced improved reasoning, multilingual support, and coding capabilities, with models scaling up to 72.71B parameters [Yang et al., 2024, Zhu et al., 2025].
@@
Besides, established in 2023 as a research initiative to push the boundaries of artificial general intelligence (AGI), DeepSeek models set out to overcome existing limitations by developing specialized models focused on efficiency, adaptability, and domain expertise [Bi et al., 2024, Dai et al., 2024, Guo et al., 2024, 2025, Liu et al., 2024b, Kasneci et al., 2023, Liu et al., 2024a, Shao et al., 2024]. In 2024, the Mixture-of-Experts (MoE), an efficiency-driven architecture that leverages sparse activation to reduce computational overhead was introduced [Guo et al., 2024, Liu et al., 2024a, b]. Meanwhile, DeepSeek Math, trained on 120B math-related tokens, was developed to handle advanced mathematical and symbolic reasoning tasks [Shao et al., 2024]. In 2025, DeepSeek made a significant breakthrough with R1 Zero, incorporating self-verification, reflection, and extended CoTs. Notably, DeepSeek-R1 was recently presented, specifically designed for mathematical, coding, and logical problem-solving, enhancing autonomous decision-making and precision in both research and enterprise applications [Guo et al., 2025, 2024, Shao et al., 2024].
@@
Notably, mathematical reasoning poses intricate challenges without straightforward solutions [Ahn et al., 2024b, Liang et al., 2024]. Unlike routine tasks that follow established frameworks, mathematical reasoning demands creativity, abstract thinking, and advanced cognitive skills. In this work, we systematically evaluate the mathematical reasoning performance of a wide array of LLMs, with a particular emphasis on DeepSeek models, to further unveil machine creativity in mathematical reasoning. Leveraging multiple independent datasets and quantitative metrics, we conduct a comprehensive empirical analysis to assess and compare their capabilities. Our findings render a detailed evaluation of prominent LLMs in mathematical reasoning while highlighting the advancements and unique strengths of LLMs.
2 Experimental Framework
@@
In this section, we provide a comprehensive evaluation of our experimental framework, detailing the datasets utilized, the quantitative evaluation metrics, as well as the overall methodology. We also present a justification for selecting these datasets as benchmarks, emphasizing their relevance, diversity, and suitability in effectively assessing LLM performance in mathematical reasoning. By establishing a consistent evaluation framework, we aim to ensure a thorough and insightful comparison of model capabilities across various reasoning tasks.
Figure 1: Overview of the experimental framework: (A) Task preparation involves selecting benchmark datasets spanning academic, graduate-level, and competition math problems. (B) Zero-shot prompting is used to query various LLMs with structured instructions. (C) Model outputs are evaluated using DeepEval across reasoning, clarity, and correctness dimensions, enabling quantitative comparison of model performance.
2.1 Experimental Designs
The framework of this empirical study is illustrated in Figure 1 and consists of three sections: A) We selected benchmarks consisting of diverse math datasets. [Cobbe et al., 2021, Hendrycks et al., 2020, Lu et al., 2023] B). Zero-Shot Prompting: Zero-shot prompting was employed to design all prompts [Hu et al., 2025], ensuring a standardized evaluation approach. The prompts are uniformly applied across a suite of language models. [Imran and Almusharraf, 2024, Achiam et al., 2023, Guo et al., 2025, Arrieta et al., 2025, Zhong et al., 2024, Liu et al., 2024a](C) We evaluate model outputs using DeepEval’s [Ip and Vongthongsri, 2025] fine-grained metrics: correctness, clarity, and reasoning, each computed by comparing predicted answers to ground truth under structured evaluation steps. Overall, this structured framework enabled holistic and systematic evaluations of LLMs in mathematical reasoning.